OpenStudy (anonymous):

how can i find the time in kinematic equation y=Vot+1/2At^2 if Vo is given,?

OpenStudy (anonymous):

U have the eq with numbers

OpenStudy (anonymous):

Vo=15m/s A=gravity what is the time when it reach the highest position. it is thrown upward

OpenStudy (anonymous):

hey u can use another formula to find t

OpenStudy (anonymous):

v = u + at

OpenStudy (anonymous):

here v=0 becoz at maximum height velocity is 0

OpenStudy (anonymous):

a = -g becoz it is thrown up!

OpenStudy (anonymous):

what is the time when it reach the highest position?

OpenStudy (anonymous):

t = u/g

OpenStudy (anonymous):

nice tnx.. ^^

OpenStudy (anonymous):

yw..00